# QGFN: Controllable Greediness with Action Values

### v1

@@ 5 Main results
We experiment on 4 standard tasks, detailed below, which were used in prior GFlowNet literature. As baselines, we use Trajectory Balance, Sub-Trajectory Balance, LSL-GFN (Kim et al., 2023) i.e. learning to scale logits which controls greediness through temperature-conditioning, as RL baselines A2C (Mnih et al., 2016) and DQN (Mnih et al., 2013) (which on its own systematically underperforms in these tasks).
@@
Fragment-based molecule generation task: Generate a graph of up to 9 fragments, where the reward is based on a prediction of the binding affinity to the sEH protein, using a model provided by Bengio et al. (2021a). $|\mathcal{X}|\approx 10^{100}$, there are 72 available fragments, some with many possible attachment points. We use Tanimoto similarity (Bender & Glen, 2004), with a threshold of 0.65, and a reward threshold of 0.97. Results are shown in Fig. 2.
@@
RNA-binding task: Generate a string of 14 nucleobases. The reward is a predicted binding affinity to the target transcription factor, provided by the ViennaRNA package for the binding landscapes (Lorenz et al., 2011); we experiment with two RNA binding tasks; L14-RNA1, and L14-RNA1+2 (two binding targets) with optima computed from Sinai et al. (2020). $|\mathcal{X}|$ is $4^{14}\approx 10^{9}$, there are 4 tokens: adenine (A), cytosine (C), guanine (G), uracil (U). Results are shown in Fig. 4.
Prepend-Append bit sequences: Generate a bit sequence of length 120 in a prepend-append MDP, where $|\mathcal{X}|$, limited to $\{0,1\}^{n}$, is $2^{120}\approx 10^{36}$. For a sequence of length $n$, $R(x)=\exp\left(1-\min_{y\in M}d(x,y)/n\right)$. A sequence is considered as a unique mode if it is within the edit distance threshold $\delta$ from $M$, where $M$ is defined as per Malkin et al. (2022a) (it is also worth noting that the task we consider here is a more complex version, introduced by Shen et al. (2023), since prepend actions induce a DAG instead of a simpler tree). In our experiment, $|M|=60,n=120,k=1,\delta=28$. Results are shown in Fig. 5.
Figure 2: Fragment-based molecule generation task. Left: Average rewards over the training trajectories. Center: Number of unique modes with a reward threshold exceeding 0.97 and pairwise Tanimoto similarity score less than 0.65. Right: Average pairwise Tanimoto similarity score for the top 1000 molecules sampled by reward. Lines are the interquartile mean and standard error calculated over 5 seeds.
Figure 4: RNA-binding tasks: Left: Average reward and modes for L14RNA1 task. Right: Average reward and modes for L14RNA1+2 task, based on 5 seeds (interquartile mean and standard error shown).
Figure 5: Bit sequence generation task. Interquartile mean and standard error of with $k=1$ over 5 seeds.
@@ 5.1 Analysis of results
Across tasks, QGFN variants produce higher rewards and find a higher number of modes, i.e. high-reward dissimilar objects. This could seem surprising, because a priori, increasing the greediness of a method likely reduces its diversity. This fundamental trade-off is known in RL as the exploration-exploitation dilemma (Sutton, 1988; Sutton & Barto, 2018). However, we are leveraging two methods and combining their strengths to reap the best from both worlds: GFlowNets are able to cover the state space, because they attempt to model all of it, by learning $p_{\theta}(x)\propto R(x)$, while $Q$ approximates the expected reward of a particular action, which can help guide the agent by selecting high-expected-reward branches. Another way to think about this: GFNs are able to estimate how many high-reward objects there are in different parts of the state space. The agent thus ends up going in all important regions of the state space, but by being a bit more greedy through $Q$, it focuses on higher reward objects, so it is more likely to find objects with reward past the mode threshold.
We also report the average reward and pairwise similarity based on 1000 samples taken after training in Table 1. Again, QGFNs outperform GFNs in reward, while retaining low levels of inter-sample similarity. We also note that at inference, we are able to use a different (and better) $p$ value than the one used at training time. We expand on this in §6, and show that it is easy to tune $p$ to achieve different reward-diversity trade-offs at inference.
Table 1: Fragment-based Molecule Generation Task: Reward and Diversity at inference with pre-trained models. The $p$ values for $p$-greedy, $p$-of-max, and $p$-quantile QGFN are set at 0.4, 0.9858, and 0.93, respectively.
@@
QGFN variants matter
We point the reader to an interesting result, which is consistent with our understanding of the proposed method. In the fragment task, the number of actions available to the agent is quite large, ranging from about 100-1000 actions, and the best performing QGFN variant is one that consistently masks most actions: $p$-quantile QGFN. It is likely indeed that most actions are harmful, as combining two fragments that do not go together may be irreversibly bad, and masking helps the agent avoid undesirable regions of the state space. However, masking a fixed ratio of actions can provide more stable training.
@@ 6 Method Analysis
We now analyze the key design choices in the methods we presented. We start by investigating the impact of $n$ (the number of bootstrapping steps in $Q$-Learning) and $p$ (the mixture parameter) on training. We then look at trained models, and reuse the learned $Q$ and $P_{F}$ to show that it is possible to use a variety of sampling strategies, and to change the mixture factor $p$ to obtain a spectrum of greediness. Finally, we run analysis experiments to provide evidence as to why QGFN is helpful.
Impact of $\beta$ in QGFN
@@
Fig. 6 shows the effect of training with different $\beta$ values on the average reward and number of modes when taking 1000 samples after training is done in the fragment task (over 5 seeds). As predicted, increasing $\beta$ increases the average reward of the agent, but at some point, causes it to become dramatically less diverse. As discussed earlier, this is typical of GFNs with a too high $\beta$, and is caused by a collapse around high-reward points and an inability for the model to further explore. While QGFN is also affected by this, it does not require as drastic values of $\beta$ to obtain a high average reward.
@@
Figure 6: Effect of $\beta$; fragment task. Increasing greediness through $\beta$ increases the average reward, but can cause diversity collapse. QGFN maintains diversity with a lower $\beta$, while GFN collapses. Modes are counted from 1000 samples at inference, using an inference-adjusted $p$.
Figure 7: Effect of training $p$, and $n$; fragment task. Changing $p$ can control greediness. Increasing $n$ is generally beneficial. Modes are counted from 1000 samples generated using the training $p$.
Impact of $n$ in QGFN
@@
As mentioned in §4, training $Q$ with 1-step returns is ineffective and produces less useful approximations of the action value. Fig. 7 shows the number of modes from 1000 post-training samples in the fragment tasks, for models trained with a variety of $n$-step values. Models start being consistently good at $n=5$ and values close to the maximum length of a trajectory tend to work well too.
Impact of the training $p$ in QGFN
@@
While our method allows changing $p$ more or less at will, we still require some value during training. Fig. 7 shows that there are clear trade-offs between choices of $p$, some yielding significantly better diversity than others. For example, $p$-of-max is fairly sensitive to the chosen value during training, and for the fragment task doesn’t seem to perform particularly well during training. On the other hand, as we will see in the next paragraph (and is also seen in Fig. 6), $p$-of-max excels at inference, and is able to generate diverse and high-reward samples by adjusting $p$.
Changing strategy after training
@@
We now look at the impact of changing the mixture parameter $p$ and the sampling strategy for already trained models on average reward and average pairwise similarity. We use the parameters of a model trained with $p$-greedy QGFN, $p=0.4$.
@@
Finally, these results suggest that we should be able to prototype new QGFN variants, including expensive ones (e.g. MCTS) without having to retrain anything. We illustrate the performance of a few other variants in §E.1, Fig. 12.
Is $Q$ calibrated?
For our intuition of why QGFN works to really pan out, $Q$ has to be accurate enough to provide useful guidance towards high-reward objects. We verify that this is the case with the following experiment. We take a trained QGFN model ($p$-greedy, $p=0.4$, fragment task, maximum $n$-step) and sample $64$ trajectories. For each of those trajectories, we take a random state within the trajectory as a starting point, thereafter generating $512$ new trajectories. We then use the reward of those $512$ trajectories as an empirical estimate $\hat{Q}$ of the expected return, which $Q$ should roughly be predicting. Fig. 9 shows that this is indeed the case. Although $Q$ is not perfect, and appears to be underestimating $\hat{Q}$, it is making reasonable predictions.
Is $Q$ really helpful?
In this experiment, we verify our intuition that pruning based on $Q$-values is helpful. We again take a trained model for the fragment task, and sample $512$ trajectories. We use $p$-of-max QGFN ($p=0.95$), and compare it to a strategy estimating best pruned actions: for each trajectory, after some random number of steps $t\sim U[4,20]$ (the max is 27), we start deterministically selecting the action that is the most probable according to $P_{F}$ but would be masked according to $Q$. To ensure that this is a valid thing to do, we also simply look at Best actions, i.e. after $t\sim U[4,20]$ steps, deterministically select the action that is the most probable according to $P_{F}$, regardless of $Q$.
@@
Figure 10: Pruning helps avoid low-reward parts of the state space. Reward distributions obtained when (a) normal sampling with $p$-of-max QGFN; (b) greedily selecting actions that $p$-of-max QGFN would prune, but are the most likely according to $P_{F}$, Best pruned actions; (c) selecting most likely $P_{F}$ actions regardless of $Q$, Best actions; and (d) normal sampling from $P_{F}$ (without using $Q$).
Why does changing $p$ work?
Recall that for QGFN to be successful, we rely on $n$-step TD, and therefore on somewhat “on-policy” estimates of $Q^{\mu}$. $\mu$ is really $\mu_{p}$, a function of $p$, meaning that if we change $p$, say to $p^{\prime}$, during inference, $Q^{\mu_{p}}$ is not an accurate estimate of $Q^{\mu_{p^{\prime}}}$. If this is the case, then there must be a reason why it is still helpful to prune based on $Q^{\mu_{p}}$ while using $\mu_{p^{\prime}}$. In Fig. 11,we perform the same measurement as in Fig. 9, but we change the $p$ value used to measure $\hat{Q}^{\mu_{p^{\prime}}}$. We find that, while the rank correlation drastically goes down (although it stays well above 0), $Q^{\mu_{p}}$ remains helpful because it lower bounds $\hat{Q}^{\mu_{p^{\prime}}}$. If we prune based on $Q^{\mu_{p}}$, then we would want it to not lead us astray, and at least make us greedier as we increase $p$. This means that if an action is not pruned, then we expect samples coming from it to be at least as good as what $Q^{\mu_{p}}(s,a)$ predicts (in expectation). This is indeed the case.
Note that reducing $p$ simply leads $\mu$ to behave more like $P_{F}$, which is still a known-to-work reward-proportional sampler, and to rely less on $Q$, whose imperfections will then have less effect anyways.
@@ 7 Conclusion and Discussion
What didn’t work     The initial stages of this project were quite different to the presented method. Instead of combining RL and GFNs into one sampling policy, we instead trained two separate agents, one GFN and one DQN. Since both are off-policy methods, the agents were able to share the training data with each other. While we were hoping that sharing “greedy” DQN data with a GFN would make it better on high-reward trajectories, we found that this was not the case; instead, the DQN agent simply slowed down the whole method–despite trying a wide variety of tricks, see §E.3.
@@ Acknowledgements
Appendix A Experiment details: Fragment-based molecule generation
In our experimental setup, we follow the exact environment specifications and implementations detailed in Malkin et al. (2022a) with the proxy model, used for evaluating molecules, provided by Bengio et al. (2021b). We set a reward threshold of 0.97 to define modes, with a pairwise Tanimoto similarity criterion of less than $0.65$. RDKit (Landrum, 2013) is used to compare pairwise Tanimoto similarity.
Training of all models was conducted using the Adam optimizer over 10,000 iterations, with a mini-batch size of 64 and a learning rate set at $1\times 10^{-4}$. The architecture of the GFlowNet models is based on a graph attention transformer (Veličković et al., 2017), with 4 layers, 128-dimensional embeddings and 2 attention heads. The reward scaling parameter $\beta$ in $R^{\beta}$ was set to 32. To follow closely the original implementation of LSL-GFN described in Kim et al. (2023), we use $\beta\sim U^{[1,64]}$, where $U$ denotes a uniform distribution. Additionally, we define a simple Multi-layer Perceptron with a hidden size of 256 as the learnable scalar-to-scalar function for the LSL-GFN. For A2C, we use a learning rate of $1\times 10^{-4}$, a training epsilon of $1\times 10^{-2}$, and an entropy regularization coefficient of $1\times 10^{-3}$. For our SubTB baseline we use SubTB(1), i.e. all trajectories are averaged with equal weight. The GFN part of QGFNs is trained with TB.
We also ran an SAC (Haarnoja et al., 2018) baseline with different $\alpha$ values of 0.5, 0.7, 0.2, along with autotuning, and a $\gamma$ value of 0.99, but we were unable to get it to discover more than 50 modes for the same amount of training iterations and mini-batch sizes.
A.1 QGFN hyperparameters:
For all variants of QGFN, we employed a grid-search approach for hyperparameter tuning, with a focus on the parameters $p$ and $n$. In the $p$-greedy QGFN variant, we selected a value of $0.4$ for $p$ from the set $\{0.2,0.4,0.6,0.8\}$, and chose an $n$ of 25 from the set $\{1,2,4,8,16,24,25,26\}$. For the $p$-of-max QGFN variant, a $p$ of 0.91 was chosen from $\{0.2,0.4,0.6,0.8,0.9,0.91,0.93,0.95\}$, with $n$ again set at 25. To ensure stability throughout the training process, we applied cosine annealing with a single-period cosine schedule over 1500 steps. An additional threshold parameter of $1\times 10^{-5}$ was applied to the condition $p\max_{i}Q(s,i)>\text{threshold}$, to prevent the initial training phase from masking actions with very small values. We also introduced a clipping of the Q-value to a minimum of 0 to prevent instability during initial training. For the $p$-quantile QGFN, a $p$ of 0.8 was selected from $\{0.6,0.7,0.8,0.9,0.95\}$, with $n=25$. For all variants of QGFN, the DQN employed had a $\tau$ of 0.95, and random action probability of $\epsilon$ was set to 0.1.
Appendix B Experiment details: QM9
In this experiment, we follow the setup described by Jain et al. (2023), but only use the HOMO-LUMO gap as a reward signal. The rewards are normalized to fall between [0, 1], although the gap proxy may range from [1,2]. As mentioned in Section 5, the modes are computed with a reward threshold of 1.10 and a pairwise Tanimoto similarity threshold of 0.70. We employ the same architecture for all models as used in the fragment-based experiments. The training models are 5,000 iterations with a mini-batch size of 64, and $\beta$ is set to 32. RDKit (Landrum, 2013) is used to compare pairwise Tanimoto similarity. We train A2C with random action probability 0.01 chosen from $\{0.1,0.01,0.001\}$ and entropy regularization coefficient 0.001 chosen from $\{0,0.1,0.01,0.001\}$.
B.1 QGFN hyperparameters:
Appendix C Experiment details: RNA-binding task
We follow the setup of Jain et al. (2022) but using the task introduced in Sinai et al. (2020). We use a sequence transformer (Vaswani et al., 2017) architecture with 4 layers, 64-dimensional embeddings, and 2 attention heads. The training for this task is 5000 iterations over mini-batch sizes of 64. The reward scaling parameter $\beta$ is set to 8 and a learning rate of $1\times 10^{-4}$ and $1\times 10^{-2}$ for $\log Z$. In this task, $\beta\sim U^{[1,16]}$ is used for LSL-GFN. Following the approach described by Sinai et al. (2020), the modes are predefined from enumerating the entire RNA landscape for L14RNA1 and L14RNA1+2 to identify local optimal through exhaustive search. ViennaRNA (Lorenz et al., 2011) is used to provide the RNA binding landscape.
C.1 QGFN hyperparameters:
Appendix D Experiment details: Bit sequence generation
The bit sequence generation task follows the same environmental setup as Malkin et al. (2022a) with $\beta$ value as 3. We generate $|M|=60$ reference sequences by randomly combining $m=15$ symbols from an initial vocabulary $H=\{00000000,11111111,11110000,00001111,00111100\}$.
Beyond the original auto-regressive generative framework in Malkin et al. (2022a), we generate sequences in a prepend-apppend fashion motivated by Shen et al. (2023). We use a Transformer (Vaswani et al., 2023) architecture for all experiments with rotary position embeddings, 3 hidden layers with hidden dimension 64 across 8 attention heads. All methods were trained for 50,000 iterations with a minibatch size of 16. For trajectory balance, we use a learning rate of of $1\times 10^{-4}$ for the policy parameters and $1\times 10^{-3}$ for $\log Z$. For SubTB, we use the same hyperparameters as in Madan et al. (2023). For LSL-GFN, we use $\beta\sim U^{[1,6]}$, where $U$ denotes a uniform distribution.
D.1 QGFN hyperparameters:
Appendix E Additional experiments and plots

### v3

@@ 5 Main results
We experiment on 5 standard tasks used in prior GFlowNet literature. As baselines, we use Trajectory Balance, Sub-Trajectory Balance, LSL-GFN (Kim et al., 2023a) i.e. learning to scale logits which controls greediness through temperature-conditioning, and as RL baselines A2C (Mnih et al., 2016), DQN (Mnih et al., 2013) (which on its own systematically underperforms in these tasks), and Tiapkin et al. (2023)’s MunDQN/GFlowNet.
@@
Fragment-based molecule generation task: [^3] We note that there exist a number of differing implementations of this task in the GFlowNet literature, we use that of https://github.com/recursionpharma/gflownet Generate a graph of up to 9 fragments, where the reward is based on a prediction of the binding affinity to the sEH protein, using a model provided by Bengio et al. (2021a). $|\mathcal{X}|\approx 10^{100}$, there are 72 available fragments, some with many possible attachment points. We use Tanimoto similarity (Bender and Glen, 2004), with a threshold of 0.65, and a reward threshold of 0.97. Results are shown in Fig. 2.
@@
Figure 4: RNA-binding tasks, Average reward and modes. Left: L14RNA1 task. Right: L14RNA1+2 task, based on 5 seeds (interquartile mean and standard error shown).
Figure 5: Bit sequence task, $k=1$. Interquartile mean and standard error over 5 seeds.
RNA-binding task: Generate a string of 14 nucleobases. The reward is a predicted binding affinity to the target transcription factor, provided by the ViennaRNA (Lorenz et al., 2011) package for the binding landscapes; we experiment with two RNA binding tasks; L14-RNA1, and L14-RNA1+2 (two binding targets) with optima computed from Sinai et al. (2020). $|\mathcal{X}|$ is $4^{14}\approx 10^{9}$, there are 4 tokens: adenine (A), cytosine (C), guanine (G), uracil (U). Results are shown in Fig. 4.
Prepend-Append bit sequences: Generate a bit sequence of length 120 in a prepend-append MDP, where $|\mathcal{X}|$, limited to $\{0,1\}^{n}$, is $2^{120}\approx 10^{36}$. For a sequence of length $n$, $R(x)=\exp\left(1-\min_{y\in M}d(x,y)/n\right)$. A sequence is considered a mode if it is within edit distance $\delta$ from $M$, where $M$ is defined as per Malkin et al. (2022a) (although the task we consider here is a more complex version, introduced by Shen et al. (2023), since prepend actions induce a DAG instead of a simpler tree). In our experiment, $|M|=60,n=120,k=1,\delta=28$, where $k$ is the bit width of actions. Results are shown in Fig. 5.
@@ 5.1 Analysis of results
Across tasks, QGFN variants produce high rewards and find a higher number of modes, i.e. high-reward dissimilar objects. The latter could seem surprising, because a priori, increasing the greediness of a method likely reduces its diversity. This fundamental trade-off is known in RL as the exploration-exploitation dilemma (Sutton, 1988; Sutton and Barto, 2018). However, we are leveraging two methods and combining their strengths to reap the best from both worlds: GFlowNets are able to cover the state space, because they attempt to model all of it, by learning $p_{\theta}(x)\propto R(x)$, while $Q$ approximates the expected reward of a particular action, which can help guide the agent by selecting high-expected-reward branches. Another way to think about this: GFNs are able to estimate how many high-reward objects there are in different parts of the state space. The agent thus ends up going in all important regions of the state space, but by being a bit more greedy through $Q$, it focuses on higher reward objects, so it is more likely to find objects with reward past the mode threshold. To further understand the performance of QGFN, we formally analyse a bandit setting, and include derivations to illustrate the general case, in Appendix §A.
We also report the average reward and pairwise similarity for the fragment task based on 1000 samples over 5 seeds taken after training in Table 1. Again, QGFNs outperform GFNs in reward, while retaining low levels of inter-sample similarity. We again note that at inference, we are able to use a different (and better) $p$ value than the one used at training time. We expand on this in §6, and show that it is easy to tune $p$ to achieve different reward-diversity trade-offs at inference. The exact $p$ values used for Table 1 are provided in Appendix §E. Also note that in LSL-GFN $\beta$ is tunable at inference, and in Table 1 we choose the $\beta$ value such that average similarity is near the 0.65 mode threshold we use (choosing a greedier $\beta$ induces a collapse in diversity).
Table 1: Fragment-based molecule task: Reward and Diversity at inference after training.
@@
QGFN variants matter    We point the reader to an interesting result, which is consistent with our understanding of the proposed method. In the fragment task, the number of actions available to the agent is quite large, ranging from about 100 to 1000 actions depending on the state, and the best performing QGFN variant is one that consistently masks most actions: $p$-quantile QGFN. It is likely indeed that most actions are harmful, as combining two fragments that do not go together may be irreversibly bad, and masking helps the agent avoid undesirable regions of the state space. However, masking a fixed ratio of actions can provide more stable training.
@@ 6 Method Analysis
We now analyze the key design choices in QGFN. We start by investigating the impact of $n$ (the number of bootstrapping steps in $Q$-Learning) and $p$ (the mixture parameter) on training. We then look at trained models, and reuse the learned $Q$ and $P_{F}$ to show that it is possible to use a variety of sampling strategies, and to change the mixture factor $p$ to obtain a spectrum of greediness at test time. Finally, we empirically probe models to provide evidence as to why QGFN is helpful.
@@
Impact of $\beta$ in QGFN    Fig. 6 shows the effect of training with different $\beta$ values on the average reward and number of modes when taking 1000 samples after training is done in the fragment task (over 5 seeds). As predicted, increasing $\beta$ increases the average reward of the agent, but at some point, causes it to become dramatically less diverse. As discussed earlier, this is typical of GFNs with a too high $\beta$, and is caused by a collapse around high-reward points and an inability for the model to further explore. While QGFN is also affected by this, it does not require as drastic values of $\beta$ to obtain a high average reward and discover a high number of modes.
@@
Figure 6: Fragment task. Left: Effect of $\beta$: Increasing greediness through $\beta$ increases the average reward but may lead to diversity collapse. QGFN maintains diversity with a lower $\beta$, while GFN collapses. Modes are counted from 1000 samples at inference, using an inference-adjusted $p$. Right: Effect of training parameters $p$, and $n$: Changing $p$ can control greediness, while increasing $n$ is generally beneficial. Modes are counted from 1000 samples generated using the training $p$.
@@
Impact of $n$ in QGFN    As mentioned in §4, training $Q$ with 1-step returns is ineffective and produces less useful approximations of the action value. Fig. 6 shows the number of modes within 1000 post-training samples in the fragment tasks, for models trained with a variety of $n$-step values. Models start being consistently good at $n=5$ and values close to the maximum length of a trajectory tend to work well too.
@@
Impact of the training $p$ in QGFN    While our method allows changing $p$ more or less at will, we still require some value during training. Fig. 6 shows that there are clear trade-offs between choices of $p$, some yielding significantly better diversity than others. For example, $p$-of-max is fairly sensitive to the chosen value during training, and for the fragment task doesn’t seem to perform particularly well during training (especially when not annealed). On the other hand, as we will see in the next paragraph (and is also seen in Fig. 6), $p$-of-max excels at inference, and is able to generate diverse and high-reward samples by adjusting $p$.
@@
Changing strategy after training    We now look at the impact of changing the mixture parameter $p$ and the sampling strategy for already trained models on average reward and average pairwise similarity. We use the parameters of a model trained with $p$-greedy QGFN, $p=0.4$.
@@
Finally, these results suggest that we should be able to prototype new QGFN variants, including expensive ones (e.g. MCTS) without having to retrain anything. We illustrate the performance of a few other variants in §B.2, Fig. 12.
Is $Q$ calibrated?    For our intuition of why QGFN works to really pan out, $Q$ has to be accurate enough to provide useful guidance towards high-reward objects. We verify that this is the case with the following experiment. We take a trained QGFN model ($p$-greedy, $p=0.4$, fragment task, maximum $n$-step) and sample $64$ trajectories. For each of those trajectories, we take a random state within the trajectory as a starting point, thereafter generating $512$ new trajectories. We then use the reward of those $512$ trajectories as an empirical estimate $\hat{Q}$ of the expected return, which $Q$ should roughly be predicting. Fig. 7 shows that this is indeed the case. Although $Q$ is not perfect, and appears to be underestimating $\hat{Q}$, it is making reasonable predictions.
Figure 8: Pruning helps avoid low-reward parts of the state space. Reward distributions when (a) sampling with $p$-of-max; (b) greedily according to $P_{F}$ selecting actions that $p$-of-max would prune, Best pruned actions; (c) selecting most likely $P_{F}$ actions regardless of $Q$, Best actions; and (d) normal sampling from $P_{F}$ (without using $Q$).
@@
Is $Q$ really helpful?    In this experiment, we verify our intuition that pruning based on $Q$-values is helpful. We again take a trained model for the fragment task, and sample $512$ trajectories. We use $p$-of-max QGFN ($p=0.95$), and compare it to a strategy estimating best pruned actions: for each trajectory, after some random number of steps $t\sim U[4,20]$ (the max is 27), we start deterministically selecting the action that is the most probable according to $P_{F}$ but would be masked according to $Q$. To ensure that this is a valid thing to do, we also simply look at Best actions, i.e. after $t\sim U[4,20]$ steps, deterministically select the action that is the most probable according to $P_{F}$, regardless of $Q$.
@@
Why does changing $p$ work?    Recall that for QGFN to be successful, we rely on $n$-step TD, and therefore on somewhat “on-policy” estimates of $Q^{\mu}$. $\mu$ is really $\mu_{p}$, a function of $p$, meaning that if we change $p$, say to $p^{\prime}$, during inference, $Q^{\mu_{p}}$ is not an accurate estimate of $Q^{\mu_{p^{\prime}}}$. If this is the case, then there must be a reason why it is still helpful to prune based on $Q^{\mu_{p}}$ while using $\mu_{p^{\prime}}$. In Fig. 10,we perform the same measurement as in Fig. 7, but we change the $p$ value used to measure $\hat{Q}^{\mu_{p^{\prime}}}$. We find that, while the rank correlation drastically goes down (although it stays well above 0), $Q^{\mu_{p}}$ remains helpful because it lower bounds $\hat{Q}^{\mu_{p^{\prime}}}$. If we prune based on $Q^{\mu_{p}}$, then we would want it to not lead us astray, and at least make us greedier as we increase $p$. This means that if an action is not pruned, then we expect samples coming from it to be at least as good as what $Q^{\mu_{p}}(s,a)$ predicts (in expectation). This is indeed the case.
Note that reducing $p$ simply leads $\mu$ to behave more like $P_{F}$, which is still a good sampler, and to rely less on $Q$, whose imperfections will then have less effect anyways.
@@ 7 Conclusion and Discussion
Our analysis suggests that at training time, QGFN works because it helps the agent to “waste” less time and capacity modeling low-reward objects, and that conversely the policy family that QGFN learns is able to sample more distinct high-reward objects given the same budget. In this sense, QGFN benefits from the advantages of both the GFlowNet and “Hard“-RL frameworks.
What didn’t work    The initial stages of this project were quite different. Instead of combining RL and GFNs into one sampling policy, we instead trained two agents, a GFN and a DQN. Since both are off-policy methods we were hoping that sharing “greedy” DQN data with a GFN would be fine and make GFN better on high-reward trajectories. This was not the case, instead, the DQN agent simply slowed down the whole method–despite trying a wide variety of tricks, see §C.
Limitations    Because we train two models, our method requires more memory and FLOPs, and consequently takes more time to train compared to TB (as shown in Table 3). QGFN is also sensitive to how well $Q$ is learned, and as we’ve shown $n$-step returns are crucial for our method to work. In addition, although the problems we tackle are non-trivial, we do not explore the parameter and compute scaling behaviors of the benchmarked methods.
@@ Acknowledgements
Appendix A Analysing $p$-greedy
Consider the bandit setting where trajectories are 1 step and just consist in choosing a terminal state. Let $p_{G}(s)=R(s)/Z$. Let $10<p<$ 1, then with $\mu(s^{\prime}|s)=(1-p)P_{F}(s^{\prime}|s)+p\mathbb{I}[s^{\prime}=\arg\max Q(s,s^{\prime})]$, assuming there is only a single argmax $s^{*}$, then $p_{\mu}(s)=(1-p)R(s)/Z+p\mathbb{I}[s=\arg\max R(s)]$. This means that for every non-argmax state, $p_{\mu}(s)=(1-p)p_{G}(s)<p_{G}(s)$. We get that $\mathbb{E}_{\mu}[R]>\mathbb{E}_{G}[R]$:
$\mathbb{E}_{\mu}[R]-\mathbb{E}_{G}[R]=$ $\sum_{s}p_{\mu}(s)R(s)-\sum_{s}p_{G}(s)R(s)$ (6)
$=$ $(p+(1-p)R(s^{*})/Z-R(s^{*})/Z)R(s^{*})+\sum_{s\neq s^{*}}(1-p)R(s)^{2}/Z-R(s)^{2}/Z$ (7)
$=$ $pR(s^{*})-pR(s^{*})^{2}/Z+\sum_{s\neq s^{*}}(-p)R(s)^{2}/Z$ (8)
$=$ $p/Z\left(R(s^{*})Z-R(s^{*})^{2}-\sum_{s\neq s^{*}}R(s)^{2}\right),\;\;\;Z=\sum_{s}R(s)$ (9)
$=$ $p/Z\left(R(s^{*})[\sum_{s}R(s)]-R(s^{*})^{2}-\sum_{s\neq s^{*}}R(s)^{2}\right)$ (10)
$=$ $p/Z\left(R(s^{*})[\sum_{s}R(s)]-\sum_{s}R(s)^{2}\right)$ (11)
$=$ $p/Z\left(\sum_{s}R(s^{*})R(s)-R(s)^{2}\right)$ (12)
since $R(s^{*})>R(s)$ and both are positive then $R(s^{*})R(s)>R(s)^{2}$ thus the last sum is positive. All other terms are positive, therefore $\mathbb{E}_{\mu}[R]-\mathbb{E}_{G}[R]>0$.
In the more general case, we are not aware of a satisfying closed form, but consider the following exercise.
Let $m(s,s^{\prime})=\mathbb{I}[s^{\prime}=\arg\max Q(s,s^{\prime})]$. Let $F^{\prime}$ be the "QGFN flow" which we’ll decompose as $F^{\prime}=F_{G}+F_{Q}$ where we think of $F_{G}$ and $F_{Q}$ as the GFN and Q-greedy contributions to the flows. Then:
$F^{\prime}(s)$ $=\sum_{z\in\mathrm{Par}(s)}F^{\prime}(z)((1-p)P_{F}(s|z)+pm(z,s))$ (14)
$=\sum_{z}F^{\prime}(z)(1-p)P_{F}(s|z)+\sum_{z}F^{\prime}(z)pm(z,s)$ (15)
$=\sum_{z}F_{G}(z)(1-p)P_{F}(s|z)+F_{Q}(z)(1-p)P_{F}(s|z)+\sum_{z}F^{\prime}(z)pm(z,s)$ (16)
$=(1-p)F_{G}(s)+\sum_{z}F_{Q}(z)\mu(z|s)+F_{G}(z)pm(z,s)$ (17)
Recall that $p(s)\propto F(s)$. Intuitively then, the probability of being in a state is reduced by a factor $(1-p)$, but possibly increased by this extra flow that has two origins. First, flow $F_{Q}$ carried over by $\mu$, and second, new flow being "stolen" from $F_{G}$ from parents when $m(z,s)=1$, i.e. when $s$ is the argmax child.
This suggests that flow (probability mass) in a $p$-greedy QGFN is smoothly redirected towards states with locally highest reward from ancestors of such states. Conversely, states which have many ancestors for which they are not the highest rewarding descendant will have their probability diminished.
Appendix B Additional experiments and analyses
In this section, we provide additional experiments to support our main findings. We explore the use of $Q$ functions from different behavior policies, assess various QGFN inference variants, examine QGFN variants trained with alternative objectives, and investigate the effects of weight sharing in QGFN models.